Cinco De Mayo has a target debt-equity ratio of 0.62. (Hint: if youre not sure what to do about debt-equity ratio, watch m
Answer #1
D/A = D/(E+D)
D/A = 0.62/(1+0.62)
=0.3827
Weight of equity = 1-D/A
Weight of equity = 1-0.3827
W(E)=0.6173
Weight of debt = D/A
Weight of debt = 0.3827
W(D)=0.3827
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 10*(1-0.31)
= 6.9
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=6.9*0.3827+19*0.6173
WACC =14.37%